a) a horizontal long straight wire carries a current of I1= 8A from left to right. The wire is tangential to a circular loop of wire of Radius R= .03 m lying immediately below it. The loop carries a current of l2= 2A in the counter clockwise direction. Assuming that the thickness of the wires are negligible, find the magnitude and direction of the net magnetic field at the center of the circular loop b) a second long wire also carrying a current l1= 8A from left to right is added and lies parallel to the first wire, tangential to the loop but immediatley below the loop. Find the magnitude and direction of the net magnetic field at the center of the circular loop. c) find the magnitude and direction of the net magnetic field at the center of the circular loop if the direction of the current I1 = 8A in the second wire lying below the loop is reversed so that it flows from right to left d) find the magnetic field at a point midway between the two horizontal wires if the circular loop is removed and each of the wires carries a current I1= 8A from left to right. The two wires are as before .03 m from each other
Answer #1

using right hand thumb rule,

direction of magnetic field due to wire due to is downward at center of circle,whereas due to loop, it is upward

so net magnetic field

B=Bw-Bl

Bw=u0*i1/2*pi*r

Bl=u0*i2/2*r

we know i1, i2 and r

so we can know net B

direction of net magnetic field, if the magnitude due to wire is more it will be downwards, otherwise it is upwards.

we can find other parts also , using same method.